# Quantum-limited superresolution of two incoherent point sources with unknown photon numbers

### v1

@@ V.2 Gaussian point-spread functions with arbitrary widths $\sigma_{1}$ and $\sigma_{2}$
Table 2 summarizes the results of $\mathcal{H}_{d}$ in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (52) for two arbitrary Gaussian point-spread functions in the limit $d\rightarrow 0$ for all feasible ranges of the exponents $t$ and $s$, which shows that different magnitudes of the difference in the photon numbers and the difference in the widths lead to different estimation precisions of the centroid separation. Similar as the general results in the preceding section, the resurgence of Rayleigh’s curse occurs only in the case that the widths of the two Gaussian point-distribution functions are sufficiently close, i.e., $t\geq 2$, and the difference of the photon numbers keeps finite compared to the total photon number, i.e., $\epsilon=O(1)$. For all the other cases, the quantum superresolution scheme can work. The case $t\geq 2$, i.e., the ratio of the difference of the widths to the total of the widths is of order $O(\widetilde{d}^{2})$ or higher, with $\epsilon\leq O(\widetilde{d}^{2})$, includes the case considered by the original superresolution scheme [20] for two identical Gaussian optical sources with equal photon numbers. For the other regimes of the above two ratios, the quantum resolution protocol is still effective, but the estimation precision may vary. In particular, when $t\leq 1$, i.e., the order of the ratio between the difference and the total of the two Gaussian widths is lower than or equal to $O(\widetilde{d})$, the superresolution can always work regardless of the photon number difference, implying the difference between the two Gaussian point-spread functions is sufficient to make the quantum superresolution work in this case, which is a manifestation of the effect of the difference between the point-spread functions in extending the effective regime of quantum superresolution.
@@
The superresolution precision $\mathcal{H}_{d}$ for two different Gaussian point-spread functions is plotted in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with respective to $\epsilon$ in Fig. 3. It can be seen from Fig. 3a that given the centroid separation $d=0$, for each $\eta$, the optimal estimation precision can always be achieved, but by different $\epsilon$. In fact, the estimation precision reaches the maximum when $\epsilon=\eta$, which can be obtained by optimizing Eq. (50) over $\epsilon$ and $\eta$ as discussed above. And the figure verifies that optimization result. Fig. 3b considers the case $\eta=0.005$ for different ratios between the centroid separation $d$ and the total width $\sigma_{{\rm tot}}$ of the two Gaussian point-spread functions. It shows that the estimation can always reach the highest precision when the photon numbers of the two sources are equal, i.e., $\epsilon=0$. And as $\epsilon$ increases, the estimation precision $\mathcal{H}_{d}$ will decrease to zero, which is the resurgence of the Rayleigh’s curse.
@@ VI Conclusions
The superresolution is an important quantum technique to overcome the Rayleigh’s curse in resolving two closely located point sources, manifesting the power of quantum estimation theory in the field of quantum imaging. Nevertheless, the resurgence of Rayleigh’s curse when the photon numbers of the two point sources are unknown imposes limitations on its applicability. This work studies the condition for the quantum superresolution technique to work by analyzing the competing roles of the difference in the photon numbers and the difference in the point-spread functions between the two optical sources as well as the centroid separation itself in determining the estimation precision of the centroid separation. The results not only provide conditions for the realizability of quantum superresolution, but also show various precision limits of superresolution in different regimes of parameters, which indicates that the competition between the difference of the photon numbers, the difference of the point-spread functions and the magnitude of the centroid separation plays a key role in the validity of quantum superresolution. We hope this work can provide new insights into the quantum superresolution theory and extend the applications of quantum superresolution in practice.
@@ Appendix A Precision limit of direct imaging
Suppose the distribution of photons on the image plane is
@@
$\epsilon=\frac{N_{2}-N_{1}}{N_{1}+N_{2}}.$ (55)
In order to quantify the precision limit of estimating the centroid separation of two incoherent optical sources, the well-known Cramér-Rao bound [15, 18, 19, 23] from the classical parameter estimation theory needs to be invoked, which demonstrates that over all possible estimation strategies, the covariance matrix $C$ for the estimates of the parameters $g_{1},\cdots,g_{n}$ is lower bounded by
@@
For the task of resolving two point sources, the centroid positions of the two sources, $X_{1}$ and $X_{2}$, or equivalently the mean and difference of the two centroid positions, $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate, and $P(x)$ is the position distribution $\Lambda(x)$ (54) of the photons from the incoherent sources.
For the direct imaging with an ideal continuum photodetector in the image plane, if the intensities and the point-spread functions of the two optical sources are equal, the position distribution (54) of the arrival photons can be rewritten as
@@
When the centroid separation $d$ is small, one can expand $\Lambda(x)$ to the second order of $d$,
@@
$\mathcal{J}=\begin{bmatrix}\alpha_{1}-\frac{\alpha_{2}d^{2}}{4}&0\\ 0&\frac{\alpha_{2}d^{2}}{16}\end{bmatrix},$ (61)
@@
${\rm Var}(\hat{d})$ $\geqslant{\rm Tr}(WJ^{-1}),$ (63)
where $\hat{d}$ denotes the estimate of the centroid separation $d$. Considering the variance may diverge when $d\rightarrow 0$, we use the reciprocal of the variance to quantify the estimation precision of $d$. Thus the estimation precision of the centroid separation $d$ by direct imaging is
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{1}{16}d^{2}\alpha_{2}N_{\text{tot}}.$ (64)
Obviously, when $d\rightarrow 0$, the estimation separation precision vanishes, so one can hardly obtain any information about the centroid separation by direct imaging when the two point sources are sufficiently close to each other. This is the Rayleigh’s curse for direct imaging of two identical optical point sources.
The above result can be generalized to the case that the two point-spread functions are the same but the photon intensities are unequal. In this case, the position distribution of the photons can be written as
@@
$\mathcal{J}=\begin{bmatrix}\alpha_{1}-\frac{\alpha_{2}\left(1-\epsilon^{2}\right)}{4}d^{2}&\frac{\alpha_{1}\epsilon}{2}+\frac{\alpha_{2}\epsilon}{16}d^{2}\\ \frac{\alpha_{1}\epsilon}{2}+\frac{\alpha_{2}\epsilon}{16}d^{2}&\frac{\alpha_{1}\epsilon^{2}}{4}+\frac{\alpha_{2}}{16}d^{2}\end{bmatrix}.$
Then the estimation precision of centroid separation by direct imaging is
@@
which means that no information about the centroid separation $d$ can be acquired by direct imaging when $d\rightarrow 0$ in this generalized case.
@@ Appendix B Multiparameter quantum metrology theory for superresolution
As the density matrix (1) of the two point sources is low-rank, it is more convenient to compute the quantum Fisher information matrix by the support of the density matrix $\hat{\rho}$,
@@
For the quantum superresolution of two point sources, the centroid positions of the two point sources, $X_{1}$ and $X_{2}$, or equivalently the mean and difference of the two centroid positions, $\bar{X}=\frac{X_{1}+X_{2}}{2}$, and $d=X_{2}-X_{1}$, are the unknown parameters to estimate, so the multi-parameter quantum Cramér-Rao bound (68) tells that
@@
where $N_{{\rm tot}}$ is the total number of photons $N_{1}+N_{2}$, $C[\bar{X},\,d]$ is the covariance matrix for the estimates of $\bar{X},\,d$, $\mathcal{Q}[\bar{X},\,d]$ is the quantum Fisher information matrix about $\bar{X},\,d$ with entries
and ^$\hat{\mathcal{L}}_{i}$ and ^$\hat{\mathcal{L}}_{j}$ are the symmetric logarithmic derivatives (SLD) of the density operator $\hat{\rho}$,
$\frac{1}{2}(\hat{\rho}\hat{\mathcal{L}}_{i}+\hat{\mathcal{L}}_{i}\hat{\rho})=\partial_{g_{i}}\hat{\rho},\;g_{i}=\bar{X},d.$ (78)
If one is interested in only one or a few unknown parameters, e.g., the centroid separation as in the superresolution, a positive-semidefinite weight matrix $W$ can be imposed on both sides of the quantum Cramér-Rao inequality (76), then the matrix quantum Cramér-Rao bound turns to a scalar form,
@@
Note that for quantum multiparameter estimation, the optimal measurements for different parameters are not always compatible, so the matrix quantum Cramér-Rao bound is not always achievable. The saturation condition [99] for the above scalar quantum Cramér-Rao bound is
@@
as the scalar quantum Cramér-Rao bound $N_{tot}^{-1}{\rm Tr}(W\mathcal{Q}^{-1})$ coincides with the Holevo bound in this case and the latter is always asymptotically achievable when $N_{{\rm tot}}\rightarrow\infty$. As the point-spread functions $\psi_{i}(x)$ are real in the quantum superresolution problem, the density matrix $\hat{\rho}$ and the eigenvalues and eigenstates of $\hat{\rho}$ are also real in this case as will be shown in the following sections, so the compatibility condition (80) is saturated, and the scalar quantum Cramér-Rao bound (79) can be achieved by the maximum-likelihood estimation strategy for $N_{{\rm tot}}\rightarrow\infty$.

### v2

@@ V.2 Gaussian point-spread functions with arbitrary widths $\sigma_{1}$ and $\sigma_{2}$
Table 2 summarizes the results of $\mathcal{H}_{d}$ in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (53) for two arbitrary Gaussian point-spread functions in the limit $d\rightarrow 0$ for all feasible ranges of the exponents $t$ and $s$, which shows that different magnitudes of the difference in the photon numbers and the difference in the widths lead to different estimation precisions of the separation. Similar as the general results in the preceding section, the Rayleigh’s criterion applies only in the case that the widths of the two Gaussian point-distribution functions are sufficiently close, i.e., $t\geq 2$, and the difference of the photon numbers keeps finite compared to the total photon number, i.e., $\epsilon=O(1)$. For all the other cases, the superresolution scheme can work. The case $t\geq 2$, i.e., the ratio of the difference of the widths to the total of the widths is of order $O(\widetilde{d}^{2})$ or higher, with $\epsilon\leq O(\widetilde{d}^{2})$, includes the case considered by the original superresolution scheme [20] for two identical Gaussian optical sources with equal photon numbers. For the other regimes of the above two ratios, the quantum resolution protocol is still effective, but the estimation precision may vary. In particular, when $t\leq 1$, i.e., the order of the ratio between the difference and the total of the two Gaussian widths is lower than or equal to $O(\widetilde{d})$, the superresolution can always work regardless of the photon number difference, implying the difference between the two Gaussian point-spread functions is sufficient to make the superresolution work in this case, which is a manifestation of the effect of the difference between the point-spread functions in extending the effective regime of superresolution.
@@
The superresolution precision $\mathcal{H}_{d}$ for two different Gaussian point-spread functions is plotted in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with respective to $\epsilon$ in Fig. 3. It can be seen from Fig. 3a that given the separation $d=0$, for each $\eta$, the optimal estimation precision can always be achieved, but by different $\epsilon$. In fact, the estimation precision reaches the maximum when $\epsilon=\eta$, which can be obtained by optimizing Eq. (51) over $\epsilon$ and $\eta$ as discussed above. And the figure verifies that optimization result. Fig. 3b considers the case $\eta=0.005$ for different dimensionless separations $\widetilde{d}$ of the two Gaussian point-spread functions. It shows that the estimation can always reach the highest precision and the quantum superresolution works when the photon numbers of the two sources are equal, i.e., $\epsilon=0$. And as $\epsilon$ increases, the estimation precision $\mathcal{H}_{d}$ will decrease to zero and the Rayleigh’s curse takes effect again.
Moreover, one can see that the precision $\mathcal{H}_{d}$ does not change monotonically with the dimensionless separation $\widetilde{d}$ and the ratio $\eta$. In fact, there exist extreme points of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$ and $\eta$, when $\epsilon$ is fixed.
Figure 3: Plot of the ratios between the estimation precision $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with respect to the ratio $\epsilon$ for two different Gaussian point-spread functions. (a) The estimation precision $\mathcal{H}_{d}$ for different ratios $\eta$ of the width difference to the total width of the two sources in the limit $d\rightarrow 0$. It can be observed that the precision $\mathcal{H}_{d}$ can still reach the maximum though the photon numbers of the two point sources are different if the point-spread functions of the two point sources are not identical, and the highest precision is achieved when $\epsilon=\eta$. The transition of $\mathcal{H}_{d}$ to the regime of Rayleigh’s curse starts when the difference of the photon numbers increases beyond the point $\epsilon=\eta$, and the larger $\eta$ is, the faster the estimation precision $\mathcal{H}_{d}$ decays with $\epsilon$. The exact effect of $\eta$ on the decay of the estimation precision is determined by Eq. (51). (b) The estimation precision $\mathcal{H}_{d}$ for different dimensionless separations $\widetilde{d}$ (55). The ratio $\eta$ (54) is $0.005$. The estimation precision can always reach the maximum when the photon numbers of the two sources are equal, i.e., $\epsilon=0$, and decays to zero as $\epsilon$ increases. The dimensionless separation $\widetilde{d}$ changes the decay speed of the estimation precision. $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ has two extreme points in this case: the maximal point is $\widetilde{d}=0$, and the minimal point is $\widetilde{d}\approx 0.2$, which can be determined by Eq. (57). So, when $0\leq\widetilde{d}\apprle 0.2$, $\mathcal{H}_{d}$ becomes smaller and decays faster with $\epsilon$ when $\widetilde{d}$ increases, and when $\widetilde{d}\apprge 0.2$, $\mathcal{H}_{d}$ behaves in the opposite way.
When $\eta\neq 0$, i.e., the widths of the two Gaussian point-spread functions are different, the extreme points of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$ can be worked out as
$\widetilde{d}_{1}=$ $0,$ (57)
$\widetilde{d}_{2}=$ $2\sqrt{\eta^{2}+1}\sqrt{w_{0}+1},$
where $w_{0}$ is the value of the principal branch $W_{0}$ of the Lambert W function at $\frac{\eta^{2}-1}{e\left(\eta^{2}+1\right)}$, i.e., $w_{0}=W_{0}\left[\frac{\eta^{2}-1}{e\left(\eta^{2}+1\right)}\right]$. When the ratio $\eta$ is small, we can expand the above second solution to $\widetilde{d}$ to the lowest order of $\eta$,
$\widetilde{d}_{2}\approx 2\sqrt{2\eta}.$ (58)
In order to determine the maximal point and the minimal point from the above two extremal solutions, we compute the second-order derivative of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$, and it turns out that
$\left.\partial_{\widetilde{d}}^{2}\mathcal{H}_{d}\right|_{\widetilde{d}=r_{1}}=-\frac{\left(1-\eta^{2}\right)\left(1-\epsilon^{2}\right)\left(-2\eta+\eta^{2}\epsilon+\epsilon\right)^{2}}{4\eta^{2}\left(\eta^{2}+1\right)^{2}\left(\eta^{2}-2\eta\epsilon+1\right)^{2}}<0,$ (59)
as $|\eta|<1$ and $|\epsilon|<1$. Therefore, $\widetilde{d}_{1}$ is the maximal point and $\widetilde{d}_{2}$ is the minimum point.
This is why $\widetilde{d}=0$ gives the maximal value in Fig. 3b. And in that figure, $\eta=0.005$, which leads to the minimal point $\widetilde{d}\approx 0.2$, so $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ decreases before $\widetilde{d}=0.2$ and increases after that point. The maximal and the minimal values of $\mathcal{H}_{d}$ can be obtained as
$\left.\mathcal{H}_{d}\right|_{\widetilde{d}=\widetilde{d}_{1}}=$ $\frac{1-\epsilon^{2}}{\eta^{2}-2\eta\epsilon+1}\mathcal{H}_{d}^{({\rm opt})},$ (60)
$\left.\mathcal{H}_{d}\right|_{\widetilde{d}=\widetilde{d}_{2}}=$ $\frac{\left(1-\epsilon^{2}\right)\left[\eta^{2}+w_{0}\left(\eta^{2}-2\eta\epsilon+1\right)+1\right]\mathcal{H}_{d}^{({\rm opt})}}{\left(1-\eta^{2}\right)^{2}\left(1-\epsilon^{2}\right)w_{0}+\left(1+\eta^{2}\right)\left(\eta^{2}-2\eta\epsilon+1\right)}.$
Interestingly, it can be verified that when $\widetilde{d}\rightarrow\pm\infty$, i.e., the separation between the two point sources is sufficiently large, the limit of $\mathcal{H}_{d}$ is
$\lim_{\widetilde{d}\rightarrow\pm\infty}=\frac{1-\epsilon^{2}}{\eta^{2}-2\eta\epsilon+1}\mathcal{H}_{d}^{({\rm opt})},$ (61)
which coincides with the maximal value of $\mathcal{H}_{d}$ at $\widetilde{d}=0$. This is why the value of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ cannot be larger than that at $\widetilde{d}=0$ when $\widetilde{d}$ goes beyond the minimal point in Fig. 3b though it increases with $\widetilde{d}$ in that regime.
When $\eta=0$, i.e., the widths of the two Gaussian point-spread functions are the same, it turns out that $\mathcal{H}_{d}$ has only one extreme point,
$\widetilde{d}_{3}=0.$ (62)
If we compute the second-order derivative of $\mathcal{H}_{d}$ with respect to $\widetilde{d}$, it shows that
$\left.\partial_{\widetilde{d}}^{2}\mathcal{H}_{d}\right|_{\widetilde{d}=\widetilde{d}_{3}}=\left(\frac{1}{\epsilon^{2}}-1\right)\mathcal{H}_{d}^{({\rm opt})}>0,$ (63)
as $|\epsilon|<1$. So, in sharp contrast to the case $\eta\neq 0$, the point $\widetilde{d}=0$ becomes the minimal point in the case $\eta=0$. This is why $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ is the smallest at $\widetilde{d}=0$ in Fig. 2.
(a) $\epsilon=0$
(b) $\epsilon=0.25$
(c) $\epsilon=0.49$
(d) $\epsilon=0.64$
Figure 4: The ratio $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ with respect to the $\eta$ and $\widetilde{d}$ for different $\epsilon$. The plots show the discontinuity of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ at the point $\eta=0$, $\widetilde{d}=0$. Whenever $\epsilon\neq 0$, the limit of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ is zero when $\widetilde{d}\rightarrow 0$ given $\eta=0$ and nonzero when $\eta\rightarrow 0$ given $\widetilde{d}=0$, so the limit $\widetilde{d}\rightarrow 0$ and the limit $\eta\rightarrow 0$ does not commute for the precision $\mathcal{H}_{d}$ in this case. When $\epsilon=0$, however, the values of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ in these two limits coincide, so the discontinuity of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ does not occur in this case.
Finally, it is worth noting that the precision $\mathcal{H}_{d}$ for the case $\sigma_{1}=\sigma_{2}$ and $\widetilde{d}\rightarrow 0$ in Fig. 2 behaves dramatically different from that for $\widetilde{d}=0$ and $\eta\rightarrow 0$ in Fig. 3a. The mechanism of this difference lies in the discontinuity of $\mathcal{H}_{d}$ at $\widetilde{d}=0$, $\eta=0$. It can be verified that
$\lim_{\widetilde{d}\rightarrow 0}\lim_{\eta\rightarrow 0}\mathcal{H}_{d}=$ $0,$ (64)
$\lim_{\eta\rightarrow 0}\lim_{\widetilde{d}\rightarrow 0}\mathcal{H}_{d}=$ $(1-\epsilon^{2})\mathcal{H}_{d}^{({\rm opt})},$
according to Eq. (47) and (51) whenever $\epsilon\neq 0$, so
$\lim_{\widetilde{d}\rightarrow 0}\lim_{\eta\rightarrow 0}\mathcal{H}_{d}\neq\lim_{\eta\rightarrow 0}\lim_{\widetilde{d}\rightarrow 0}\mathcal{H}_{d}.$ (65)
In Fig. 4, the precision $\mathcal{H}_{d}$ is plotted with respect to $\widetilde{d}$ and $\eta$ for different $\epsilon$ to illustrate this discontinuity. Hence, $\mathcal{H}_{d}$ has different behaviors with respect to $\epsilon$ in the above two cases.
@@ VI Conclusions
The superresolution is an important quantum technique to surpass the Rayleigh’s criterion in resolving two closely located point sources, manifesting the power of quantum estimation theory in the field of quantum imaging. Nevertheless, the vanishing of the estimation precision when the photon numbers of the two point sources are unknown imposes limitations on its applicability. This work studies the condition for the superresolution technique to work by analyzing the competing roles of the difference in the photon numbers and the difference in the point-spread functions between the two optical sources as well as the separation itself in determining the estimation precision of the separation. The results not only provide conditions for the realizability of quantum superresolution, but also show various precision limits of superresolution in different regimes of parameters, which indicates that the competition between the difference of the photon numbers, the difference of the point-spread functions and the magnitude of the separation plays a key role in the validity of superresolution. We hope this work can provide new insights into the quantum theory of superresolution and extend the applications of quantum superresolution in practice.
@@ Appendix A Precision limit of direct imaging
Suppose the probability distribution of a single photon of position $x$ on the image plane is
@@
$\epsilon=\frac{N_{2}-N_{1}}{N_{2}+N_{1}}.$ (67)
In order to quantify the precision limit of estimating the separation of two incoherent optical sources, the well-known Cramér-Rao bound [15, 18, 19, 23] from the classical parameter estimation theory needs to be invoked, which demonstrates that over all possible estimation strategies, the covariance matrix $C$ for the estimates of the parameters $g_{1},\cdots,g_{n}$ is lower bounded by
@@
For the task of resolving two point sources, the positions of the two sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate, and $P(x)$ is the position distribution $\Lambda(x)$ (66) of the photon from the incoherent sources.
For the direct imaging with an ideal continuum photodetector in the image plane, if the intensities and the point-spread functions of the two optical sources are equal, the position distribution (66) of the arrival single photon can be rewritten as
@@
When the separation $d$ is small, one can expand $\Lambda(x)$ to the second order of $d$,
@@
$\mathcal{J}=\begin{bmatrix}\alpha_{1}-\frac{\alpha_{2}d^{2}}{8}&0\\ 0&\frac{\alpha_{2}d^{2}}{16}\end{bmatrix},$ (73)
@@
${\rm Var}(\hat{d})$ $\geqslant{\rm Tr}(W\mathcal{J}^{-1}),$ (75)
where $\hat{d}$ denotes the estimate of the separation $d$. Considering the variance may diverge when $d\rightarrow 0$, we use the reciprocal of the variance to quantify the estimation precision of $d$. Thus the estimation precision of the separation $d$ by direct imaging is
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{1}{8}d^{2}\alpha_{2}\text{{\hbox{N_{}}}}.$_ . (76)
Obviously, when $d\rightarrow 0$, the estimation separation precision vanishes, so one can hardly obtain any information about the separation by direct imaging when the two point sources are sufficiently close to each other. This is called the Rayleigh’s curse [20] for direct imaging of two identical optical point sources.
The above result can be generalized to the case that the two point-spread functions are the same but the intensities are unequal. In this case, the position distribution of single photon can be written as
@@
$\mathcal{J}=\begin{bmatrix}\alpha_{1}+\frac{1}{8}d^{2}\left[2\beta\epsilon^{2}-\alpha_{2}\left(2\epsilon^{2}+1\right)\right]&\frac{\alpha_{1}\epsilon}{2}+\frac{\epsilon d^{2}}{16}\left[2\beta\epsilon^{2}-\alpha_{2}\left(2\epsilon^{2}+1\right)\right]\\ \frac{\epsilon\alpha_{1}}{2}+\frac{\epsilon d^{2}}{16}\left[2\beta\epsilon^{2}-\alpha_{2}\left(2\epsilon^{2}+1\right)\right]&\frac{\alpha_{1}\epsilon^{2}}{4}+\frac{d^{2}}{32}\left[\alpha_{2}\left(2-5\epsilon^{2}\right)+2\beta\epsilon^{4}\right]\end{bmatrix},$ (79)
[…]
@@
which means that no information about the separation $d$ can be acquired by direct imaging when $d\rightarrow 0$ in this generalized case.
@@ Appendix B Multiparameter quantum metrology theory for superresolution
As the density matrix (1) of a single photon is low-rank, it is more convenient to compute the quantum Fisher information matrix on the support of $\hat{\rho}$,
@@
For the superresolution of two point sources, the positions of the two point sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$, are the unknown parameters to estimate, so the multi-parameter quantum Cramér-Rao bound (81) tells that
@@
where $N_{{\rm tot}}$ is the total number of photons $N_{1}+N_{2}$, $C[\bar{X},\,d]$ is the covariance matrix for the estimates of $\bar{X},\,d$, $\mathcal{Q}[\bar{X},\,d]$ is the quantum Fisher information matrix about $\bar{X},\,d$ with entries in Eq. (82).
If one is interested in only one or a few unknown parameters, e.g., the separation as in the superresolution, a positive semidefinite weight matrix $W$ can be imposed on both sides of the quantum Cramér-Rao inequality (89), then the matrix quantum Cramér-Rao bound turns to a scalar form,
@@
Note that for quantum multiparameter estimation, the optimal measurements for different parameters are not always compatible, so the matrix quantum Cramér-Rao bound is not always achievable. The saturation condition given by Ragy et al. [100] for the above scalar quantum Cramér-Rao bound is
@@
as the scalar quantum Cramér-Rao bound $N_{tot}^{-1}{\rm Tr}(W\mathcal{Q}^{-1})$ coincides with the Holevo bound in this case and the latter is always asymptotically achievable when $N_{{\rm tot}}\rightarrow\infty$. As the point-spread functions $\psi_{i}(x)$ are real in the superresolution problem, the density matrix $\hat{\rho}$ and the eigenvalues and eigenstates of $\hat{\rho}$ are also real in this case as will be shown in the following sections, so the compatibility condition (91) is saturated, and the scalar quantum Cramér-Rao bound (90) can be achieved by the maximum-likelihood estimation strategy for $N_{{\rm tot}}\rightarrow\infty$.